xy-[yz-zx-{yx-(3y-xz)-(xy-zy)}]
1
Expert's answer
2013-09-25T08:14:09-0400
xy-[yz-zx-{yx-(3y-xz)-(xy-zy)}]=xy-[yz-zx-{yx-3y+xz-xy+zy}]=xy-[yz-zx-yx+3y-xz+xy-zy]=xy-yz+zx+yx-3y+xz-xy+zy=-yz+zx+yx-3y+xz+zy

If commutativity holds (i.e. xy=yx) the last expression can be simplified: 2zx+yx-3y (for real numbers it is true).